OpenStudy (anonymous):

For the following system, use the second equation to make a substitution for y in the first equation. 2x + y = 6 y = 3x + 4 What is the resulting equation? y + 2x + y = 6 2x + y + 3x + 4 = 6 2x + (3x + 4) = 6

OpenStudy (anonymous):

2x+(3x+4)=6. Option 3

OpenStudy (anonymous):

substituting second equation in one gives 5x+4=6 =>x=2/5 and y=26/5

OpenStudy (anonymous):

2x+(3x=4)=6. as simamura said